use s=n^2 to find the sum of 1+3+5+....+501.
------

Answer by stanbon(75887)

find the sum of 1+3+5+....+501
-----------------
1st find "n" if a(1) = 1 and d = 2
----
Using a(n) = a(1) + (n-1)d you get:
501 = 1 + (n-1)2
500 = 2(n-1)
250 = n-1
n = 251
----------------------
Find the sum using S(n)=(n/2)(1 + 501)
---
S(251) = 251/2(502) = 251*251 = 63001
=========================================
cheers,
Stan H.